In a particular solution, acetic acid is 12% ionized at 25°C. Calculate the pH of the solution and the mass of acetic acid (Ka=1.8x10^-5) dissolved to yield 1.00 L of solution.

pH = ?

Mass =? g
Answer #1

First we express the equation of the reaction that occurs:

HA + H2O = H3O + + A-

where in the balance:

n HA = No - N

n H3O + = N

n A- = N

according to the degree of dissociation:

N / No = 0.12

we cleared:

No = N / 0.12

We determine the equation of Ka:

Ka = nA- * nH3O + / nHA = N ^ 2 / (N / 0.12) - N

we simplify:

Ka = N / 7.33

we cleared:

N = 7.33 * Ka = 7.33 * 1.8x10 ^ -5 = 1.32x10 ^ -4

We calculate the mass:

mass = 1.32x10 ^ -4 * (60.05 g / 1 mol) = 0.008 g

We calculate the pH, of 1 L of solution:

pH = -Log [H3O +] = - Log (1.32x10 ^ -4) = 3.88